# Cartesian and spherical multipole expansions in anisotropic media

### v2

@@ III Illustration on an inverse problem
On top of appearing in natural materials (e.g., calcite [41]), electromagnetic anisotropic media are of prime importance for metamaterials (e.g., through effective medium theory [42]) which are seeing an increase of adoption from optical [43] to radio frequencies [44]. High anisotropy can be achieved through 3D printing of carbon fiber composite filament [45].
@@
$\mathbf{E}(\mathbf{r})=-i\omega\mu\iiint\text{d}^{3}\mathbf{r}^{\prime}\mathbf{G}(\mathbf{r}-\mathbf{r}^{\prime})\mathbf{J}(\mathbf{r}^{\prime})$ (43)
@@
$f(\omega)=\int\text{d}tf(t)e^{-i\omega t}$ (44)
@@
Figure 1: Flowchart of the inverse problem for $\varepsilon_{zz}=4$ and $\alpha=(0,0,2)$. The ground truth data consists of all components (pictured: $x$) of the electric field radiated by the source $\mathbf{J}(\mathbf{r})=\nabla^{\alpha}\delta(\mathbf{r})\mathbf{e}_{x}$ simulated using Meep and sampled on a sphere surrounding the source. The explaining data consists of the analytical spherical (bottom) and Cartesian (top) multipole expansion basis functions. These basis functions are multiplied by coefficients determined by an optimization algorithm, whose goal is to minimize the $L^{2}$-norm of the error between the ground truth and the multipole expansion. This error is normalized to the ground truth $L^{2}$-norm. Its complement after convergence represents the proportion of ground truth energy explained by the multipole expansion, indicated in the lower and upper corners.
@@
Figure 2: Effect of the truncation order $\ell_{\text{max}}$ on the representation power of both multipole expansions for both source types. The number of degrees of freedom (i.e., the number of parameters determined by the inverse problem) is also given on the right-hand side axis. The results circled in red are illustrated in Figure 1.
Figure 3: Moments recovered by the inverse problem for $\varepsilon_{zz}=4$ and $\ell_{\text{max}}=4$. The dot size is proportional to the maximum field amplitude of the corresponding moment.
Figure 4: Effect of the degree of anisotropy $\varepsilon_{zz}$ on the representation power of the spherical multipole expansion for the source $\mathbf{J}(\mathbf{r})=\nabla^{(0,0,2)}\delta(\mathbf{r})\mathbf{e}_{x}=\partial^{2}\delta(\mathbf{r})/\partial z^{2}\mathbf{e}_{x}$. For $\varepsilon_{zz}>10$, the analytical Green’s function was used as ground truth data.
@@
Solving this ordinary differential equation, we obtain (see Section II of the Supplementary Information)
$h(\mathbf{r})=c_{+}e^{ik_{0}z}+c_{-}e^{-ik_{0}z}+\frac{1}{2k_{0}}\delta(x,y)\text{sign}(z)\sin(k_{0}z)$ (51)
@@
Finally, one might argue that it is always possible to represent fields on a sphere by a projection onto spherical harmonics. While this is mathematically true, it is usually not helpful in physical problems. Indeed, the moments recovered by the projection do not correspond to those of the source responsible for the radiation. To illustrate, we project the $xx$ component of Green’s function on the sphere described above in isotropic ($\varepsilon_{zz}=1$) and anisotropic ($\varepsilon_{zz}=4$) media.
Figure 5: Representation power of the spherical harmonics projected onto the $xx$ component of Green’s function in isotropic and anisotropic media.
@@
It would also be possible to repeat the comparison above by projecting the fields onto the restriction on the sphere of an arbitrary basis of polynomials of three variables. Since the spherical harmonics are formed by a restriction on the unit sphere of harmonic and homogeneous polynomials of three variables, this would consist of a Cartesian counterpart to the spherical harmonics. However, such a Cartesian projection would still be blind to the propagation anisotropy. Moreover, the added degrees of freedom would also be redundant, as it has been long shown that the spherical harmonics span all square-integrable functions on the sphere [38].
@@ IV Conclusion
Furthermore, by approaching the equivalence problem from the singular behavior of the solutions, we have shown how the Cartesian multipole expansion—obtained straightforwardly from Green’s function—can be used to obtain the spherical one, which is not readily available in anisotropic media. This matters for cases where the dominant source feature is harmonic and, therefore, well-represented by the spherical approach. Also, the Laplacian-involving mapping from Cartesian to spherical approaches presented in Equation (24) can be applied to many wave-like systems where the multipole expansion is already applied. Anytime the propagation is anisotropic, the Cartesian approach is expected to offer a better representation power and, ultimately, an improved understanding of the underlying physics. The potential for improved computational methods in radiation problems, as demonstrated in our study, opens new possibilities for the application of Cartesian multipole expansions in a variety of physical systems, reflecting the critical importance of accommodating anisotropic properties in contemporary research and technology.
Notably, the discrepancy between both approaches only appears for sources of order above one; this, together with the relatively recent rise of anisotropic media such as metamaterials, might explain why the issue has not been researched until now. In addition to the electromagnetic application to optical and radio-frequency systems, our results are relevant to describing anisotropic interactions in plasma physics as discussed in Section I. Similarly, the implications for cosmological studies on anisotropic baryon acoustic oscillations and the cosmic neutrino background also highlight key examples where our findings can have an impact.
@@ Acknowledgments
Elias Le Boudec: conceptualization, methodology, software, validation, formal analysis, investigation, writing – original draft, visualization; Toma Oregel-Chaumont: investigation, writing – original draft, writing – review & editing; Farhad Rachidi: writing – review & editing, supervision, project administration, funding acquisition; Marcos Rubinstein: supervision; Felix Vega: supervision, project administration, funding acquisition

### v3

@@ III Illustration on an inverse problem
On top of appearing in natural materials (e.g., calcite [42]), electromagnetic anisotropic media commonly describe metamaterials (e.g., through effective medium theory [43]), which are seeing an increase of adoption from optical [44] to radio frequencies [45]. High anisotropy can be achieved through 3D printing of carbon fiber composite filament [46].
@@
$\mathbf{E}(\mathbf{r})=-\text{i}\omega\mu\iiint\text{d}^{3}\mathbf{r}^{\prime}\mathbf{G}(\mathbf{r}-\mathbf{r}^{\prime})\mathbf{J}(\mathbf{r}^{\prime})$ (43)
@@
$f(\omega)=\int\text{d}tf(t)\text{e}^{-\text{i}\omega t}$ (44)
@@
Figure 1: Flowchart of the inverse problem for $\varepsilon_{zz}=4$ and $\alpha=(0,0,2)$. (a) The ground truth data consists of all components (pictured: norm of the $x$ component, max-normalized; warm colors indicate high values) of the electric field radiated by the source $\mathbf{J}(\mathbf{r})=\nabla^{\alpha}\delta(\mathbf{r})\mathbf{e}_{x}$ simulated using Meep and sampled on a sphere surrounding the source. The explaining data consists of the analytical Cartesian (b) and spherical (c) multipole expansion basis functions. These basis functions are multiplied by coefficients (d) determined by an optimization algorithm (e), whose goal is to minimize the $L^{2}$-norm of the error (f) between the ground truth and the multipole expansion (g). This error is normalized to the ground truth $L^{2}$-norm. Its complement after convergence represents the proportion of ground truth energy explained by the Cartesian (h) and spherical (i) multipole expansions. The fields obtained after convergence of the minimizer are shown in (j) for the Cartesian and (k) for the spherical multipole expansion.
@@
Figure 2: Effect of the truncation order $\ell_{\text{max}}$ on the representation power of both multipole expansions (spherical: solid and solid-dotted; Cartesian: dashed) for both source types ($\alpha=(0,0,0)$: no markers; $\alpha=(0,0,2)$: markers). The number of degrees of freedom (i.e., the number of parameters determined by the inverse problem) is also given on the right-hand side axis. The results circled in red are illustrated in Figure 1.
Figure 3: Moments recovered by the inverse problem for $\varepsilon_{zz}=4$ and $\ell_{\text{max}}=4$. (a, c) Spherical multipole expansions for (a) $\alpha=(0,0,0)$ and (c) $\alpha=(0,0,2)$. (b, d) Cartesian multipole expansions for (b) $\alpha=(0,0,0)$ and (d) $\alpha=(0,0,2)$. The dot size is proportional to the maximum field amplitude of the corresponding moment. Warm colors indicate higher orders.
Figure 4: Effect of the degree of anisotropy $\varepsilon_{zz}$ on the representation power of the spherical multipole expansion for the source $\mathbf{J}(\mathbf{r})=\nabla^{(0,0,2)}\delta(\mathbf{r})\mathbf{e}_{x}=\partial^{2}\delta(\mathbf{r})/\partial z^{2}\mathbf{e}_{x}$. Solid yellow line: $\ell_{\text{max}}=2$; solid-dashed orange line: $\ell_{\text{max}}=4$; dashed red line: $\ell_{\text{max}}=6$. For $\varepsilon_{zz}>10$, the analytical Green’s function was used as ground truth data.
@@
Solving this ordinary differential equation, we obtain (see Supplementary Method 2)
$h(\mathbf{r})=c_{+}\text{e}^{\text{i}k_{0}z}+c_{-}\text{e}^{-\text{i}k_{0}z}+\frac{1}{2k_{0}}\delta(x,y)\text{sign}(z)\sin(k_{0}z)$ (51)
@@
Finally, one might argue that it is always possible to represent fields on a sphere by a projection onto spherical harmonics. While this is mathematically true, it is usually not helpful in physical problems. Indeed, the moments recovered by the projection do not correspond to those of the source responsible for the radiation. To illustrate, we project the $xx$ component of Green’s function on the sphere described at the beginning of this section in isotropic ($\varepsilon_{zz}=1$) and anisotropic ($\varepsilon_{zz}=4$) media.
Figure 5: Representation power of the spherical harmonics projected onto the $xx$ component of Green’s function in isotropic ($\varepsilon_{zz}=1$, solid blue line) and anisotropic media ($\varepsilon_{zz}=4$, dashed red line).
@@
It would also be possible to repeat this comparison by projecting the fields onto the restriction on the sphere of an arbitrary basis of polynomials of three variables. Since the spherical harmonics are formed by a restriction on the unit sphere of harmonic and homogeneous polynomials of three variables, this would consist of a Cartesian counterpart to the spherical harmonics. However, such a Cartesian projection would still be blind to the propagation anisotropy. Moreover, the added degrees of freedom would also be redundant, as it has been long shown that the spherical harmonics span all square-integrable functions on the sphere [39].
@@ IV Conclusion
Furthermore, by approaching the equivalence problem from the singular behavior of the solutions, we have shown how the Cartesian multipole expansion—obtained straightforwardly from Green’s function—can be used to obtain the spherical one, which is not readily available in anisotropic media. This matters for cases where the dominant source feature is harmonic and, therefore, well-represented by the spherical approach. Also, the Laplacian-involving mapping from Cartesian to spherical approaches presented in Equation (24) can be applied to many wave-like systems where the multipole expansion is already applied. Anytime the propagation is anisotropic, the Cartesian approach is expected to offer a better representation power and, ultimately, an improved understanding of the underlying physics.
Note that the discrepancy between both approaches only appears for sources of order above one; this, together with the relatively recent rise of manufactured anisotropic media such as metamaterials, may explain the lack of prior research. In addition to the electromagnetic application to optical and radio-frequency systems, our results are relevant to describing anisotropic interactions in plasma physics as discussed in the introduction. Similarly, the implications for cosmological studies on anisotropic baryon acoustic oscillations and the cosmic neutrino background also highlight key examples where our findings can have an impact.
@@ Acknowledgments
Elias Le Boudec: conceptualization, methodology, software, validation, formal analysis, investigation, writing – original draft, visualization; Toma Oregel-Chaumont: investigation, writing – original draft, writing – review & editing; Farhad Rachidi: writing – review & editing, supervision, project administration, funding acquisition; Marcos Rubinstein: supervision; Felix Vega: supervision, project administration, funding acquisition.
@@
Data availability statement
Source data for all figures are provided with the paper in the Supplementary Data 1.